OpenStudy (anonymous):

can anyone explain how to do this? determine all of the real number solutions. x-sqrtx=20

OpenStudy (anonymous):

Use u-substitution: let: \[u=\sqrt{x}\] Rewrite: \[u^2-u-20=0\]

OpenStudy (anonymous):

Factor

OpenStudy (anonymous):

why u\[u^{2?}\]

OpenStudy (anonymous):

You will see, factor that equation above

OpenStudy (anonymous):

\[u^{2}-u-20=0\]

OpenStudy (anonymous):

yes, factor it

OpenStudy (anonymous):

\[u^{2}-u=20 ?\]

OpenStudy (anonymous):

u=20

OpenStudy (anonymous):

\[(u-5)(u+4)=0\]

OpenStudy (anonymous):

right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ok now that we know this, we can substitute sqrt(x) back in for u

OpenStudy (anonymous):

\[u-5=0, \sqrt{x}-5=0, \sqrt{x}=5\] \[(\sqrt{x})^2=5^2, x=25\]

OpenStudy (anonymous):

16 is not a solution

OpenStudy (anonymous):

there are two roots one real one complex this cannot be

OpenStudy (anonymous):

There is only one root.

OpenStudy (anonymous):

I was able to find only 1 root which is 25 correct

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Thanks

OpenStudy (anonymous):

np

OpenStudy (anonymous):

\[u+4=0\] \[\sqrt{x}+4=0\] \[\sqrt{x}=-4\] we know that \[if \sqrt{x }\qquad x \ge 0 \qquad always\]

OpenStudy (anonymous):

Yes but you can't simply square both sides here, check 16 in the original equation, it doesn't work

OpenStudy (anonymous):

No solutions exist for \[\sqrt{x}=-4\]

OpenStudy (anonymous):

I know what you mean.. I can check it in original eq.